Adult tickets to the Snow Flake Museum cost $7 and children's tickets cost $4 each. Last week the museum sold 78 tickets and took in $387 from the sales. How many of each type of ticket did they sell?
Answer by mananth(16949)

let number of adult tickets sold be x at $7
let number of children's tickets sold be y at $4
..
x+y= 78...............1
7x+4y=387.............2
..
multiply 1 by 7 and subtract 2
7x+7y-7x-4y=546-387
3y= 159
y= 159/3
y=53 number ofchildren's tickets
adults tickets will be 78-53= 25 numbers